Find the possible value or values of r in the quadratic equation r^2 – 7r – 8 = 0.
Answer by dabanfield(803)

Factor r^2 - 7r -8 = (r-8)* (r+1)
So (r-8) * (r+1) = 0
The two solutions are r-8 = 0 and r+1 = 0
r=8 and r=-1